solve using any method:
x^2-1.3x=-0.4



Answer by edjones(8007)

x^2-1.3x=-0.4
10x^2-13x+4=0 multiply each side by 10.
10*4=40, sum of 2 factors of 40 that equal -13 is -8 and -5
10x^2-8x-5x+4=0
2x(5x-4)-(5x-4)=0 factoring by grouping.
(2x-1)(5x-4)=0
x=.5, x=.8
.
Ed